OpenStudy (texaschic101):

Can someone please help me with this....can't figure out how to set it up :)

OpenStudy (texaschic101):

In the Newham hockey league a team scores 5 points for a win, 2 for a draw and 0 if they lose. the club has played 20 games in the league and has scored 21 points. what is the smallest possible number of games the team could have lost

OpenStudy (anonymous):

woa im a little confused on this problem......one sec :/

OpenStudy (texaschic101):

I started with this.. 5w + 2d + 0l = 21 but then I do not know what to do....

OpenStudy (texaschic101):

w + d + l = 20 ??

OpenStudy (texaschic101):

should I just start subbing in numbers for l, starting with 1 and going up....trial and error ?

ganeshie8 (ganeshie8):

I think smallest number of losses is 11

OpenStudy (texaschic101):

11 loses. leaving 9 games. 1 win = 5 points 8 draws = 16 points 20 games at 21 points that works

OpenStudy (texaschic101):

is there any other way to do this except for trial and error ?

ganeshie8 (ganeshie8):

5w + 2d = 21 w + d + l = 20

ganeshie8 (ganeshie8):

from the first equation, solve \(d\) and substitute it in second equation

OpenStudy (texaschic101):

d = -5/2w + 21/2 w + (-5/2w + 21/2) + l = 20 w - 5/2w + 21/2 + l = 20 2/2w - 5/2w = 20 - 21/2 + l wait...where do I put the l

ganeshie8 (ganeshie8):

get rid of fractions first, they look clumsy

OpenStudy (texaschic101):

2/2w - 5/2w = 20 - 21/2 + l --->(2) 2w - 5w = 40 - 21 + l

ganeshie8 (ganeshie8):

w + d + l = 20 2w + 2d + 2l = 40 2w + 21 - 5w + 2l = 40 -3w + 2l = 19 2l = 19 + 3w

ganeshie8 (ganeshie8):

we want minimum value for Losses, so put w=0, that gives u minimum value for Losses. but can u really put w=0 ?

ganeshie8 (ganeshie8):

Nope, we cannot put w = 0, cuz, that will give fraction for Losses. so, put w = 1

OpenStudy (texaschic101):

21 - 19 = 3w 3 = 3w 1 = w

ganeshie8 (ganeshie8):

2l = 19 + 3*1 2l = 22 l = 11

ganeshie8 (ganeshie8):

its 2L, not 21 :|

OpenStudy (texaschic101):

oh....I see now. I was trying to use elimination...couldn't get it to work. Your a genius

ganeshie8 (ganeshie8):

\(w + d + l = 20\) \(2w + 2d + 2l = 40\) \(2w + 21 - 5w + 2l = 40\) \(-3w + 2l = 19\) \(2l = 19 + 3w\)

ganeshie8 (ganeshie8):

now see, the \(l\) and \(1\) look same when we dont use latex lol

ganeshie8 (ganeshie8):

no, you're a genius :)

OpenStudy (texaschic101):

I see that....however, I am not too good with Latex....lol

OpenStudy (texaschic101):

thank you so much for your help :)

ganeshie8 (ganeshie8):

np :)